OpenStudy (anonymous):

A fence is to be built to enclose a rectangular area of 220 square feet. The fence along three sides is to be made of material that costs 6 dollars per foot, and the material for the fourth side costs 16 dollars per foot. Find the length L and width W (with W≤L) of the enclosure that is most economical to construct.

OpenStudy (anonymous):

a and b are lengths A = 220 = a * b so a = 220/b cost = 2(6a) + 1 (6b) + 1(13b) cost = c = 12 a + 19 b c = 12(220/b)+19 b dc/db = 0 at max or min = 12(-220/b^2) + 19 220(12) = 19 b^2 solve for b and a

OpenStudy (anonymous):

HOPE THIS HELPS :)

OpenStudy (anonymous):

a and b are lengths A = 220 = a * b so a = 220/b cost = 2(6a) + 1 (6b) + 1(16b) cost = c = 12 a + 22 b c = 12(220/b)+22 b dc/db = 0 at max or min = 12(-220/b^2) + 22 220(12) = 22 b^2 solve for b and a

OpenStudy (anonymous):

SORRY THIS THE CORRECT AWNSER